# A Constant Factor Approximation for Directed Feedback Vertex Set in Graphs of Bounded Genus

### v3

@@ Appendix A Statement and proofs of topological results we use
([24] pg 308) In any minimal genus embedding of a graph, each face is homeomorphic to an open disk.
@@
- Suppose, without loss of generality, that $G_{L}$ is empty. Consider the face of $G$ that contains $C$ and intersects the left of $C$. By Youngs’s theorem, this face must be homeomorphic to a disc. But, $C$ is not contractible (else it would separate the surface $Q$ into two components). Hence, a small leftward shift of $C$ which will lie in the face $f$ will also not be contractible. This contradicts that $f$ is homeomorphic to a disc, since a disc is simply connected.
@@
- We first show that if a face contains vertices of $G_{L}$ and of $C$ (resp., of $G_{R}$ and of $C$), then it cannot contain vertices of $G_{R}$ (resp, of $G_{L}$). Suppose, for the sake of contradiction, that there is such a face $f$ of degree $d$. Let $\partial f=v_{0}\dotsm v_{d-1}$ be the boundary cycle of $f$, where $i\in\mathbb{Z}/d\mathbb{Z}$. Without loss of generality, assume that $v_{0}\in\hat{L}$ and for some $q$ $v_{1},v_{2},..,v_{q}\in C$, and $v_{q+1}\in\hat{R}$. There is are points $p_{l}$ on the edge $v_{0}v_{1}$ in the interior of $L$ and $p_{R}$ on the edge $v_{q}v_{q+r}$ in the interior of $R$. Let $h:[0,1]\rightarrow Q$ be a non-self-intersecting curve in $f$ from $p_{R}$ to $p_{L}$. Let $r_{L},r_{R}>0$ be such that $B_{Q}(p_{l},r_{L})\subset L,\ \ B_{Q}(p_{R},r_{R})\subset R$. Let $h^{L},h^{R}$ be non-self-intersecting curves in $B_{Q}(p_{l},r_{L})$ and $B_{Q}(p_{R},r_{R})$ respectively not intersecting $h$. Then $h\cup h^{L}\cup h^{R}$ satisfies the conditions of 4. Thus $h\cup h^{L}\cup h^{R}\cup g(p_{l}v_{0},v_{1},...,v_{q+1}p_{R})$ does not bound a region of the closure of $f$ homeomorphic to a disk which cannot happen because $f$ is homeomorphic to an open disk.
- Next, we show that if there is a face of $G$ that contains vertices of $G_{L}$ and of $G_{R}$ (but not of $C$), then we again arrive at a contradiction: the boundary of such a face $f$ will necessarily contain more than one connected component, since there can be no edge between vertices of $G_{L}$ and of $G_{R}$. But, this is not possible since $f$ is homeomorphic to a disc by Youngs’s theorem.
@@
It is well known (see for instance [1] page 15) that smooth surfaces $Q$ have the property that for each $v\in Q$ there is an open ball $B_{Q}(v,r_{0})$ of some small radius $r_{0}>0$ in $Q$ and a diffeomorphism $\psi$ from $B_{Q}(v,r_{0})$ to the open disk $B_{\mathbb{R}^{2}}(0,r_{0})$ of radius $r_{0}$ about the origin in the two-dimensional plane such that $\psi(v)=(0,0)$ and $\psi$ preserves distances from $v$, that is $\operatorname{dist}_{Q}(v,x)=\|\psi(v)-\psi(x)\|$, where $\operatorname{dist}_{Q}(v,x)$ is the geodesic distance from $v$ to $x$ in $Q$. For $p\in Q$, $r>0$, denote by $B(p,r)$ the open ball of radius $r$ about $p$. We now formally state and prove what 3 and 4 informally say.
@@
Lastly let $h:[0,1]\rightarrow Q$ be any curve that reaches $C^{\prime}$ from the right at a point $c_{2}=h(1)$ on $C^{\prime}$, leaves $C^{\prime}$ from the left at $c_{1}=h(0)$, that is the curve $\bar{\psi}(t)=h(1-t)$ reaches $C^{\prime}$ at $c_{1}$ from the left and $h$ is otherwise disjoint from $C^{\prime}$. Assume $c_{1}\neq c_{2}$ and let $C^{\prime}_{c_{1},c_{2}}$ be a subcurve of $C^{\prime}$ with endpoints $c_{1}$ and $c_{2}$. Then there is a curve $\hat{h}:[0,2]\rightarrow Q$ that reaches $C^{\prime}$ from the right at $c_{1}=\hat{h}(1)$ and leaves $C^{\prime}$ at a point $c_{1}=\hat{h}(0)$ $-^\hat{h}$ is otherwise disjoint from $C^{\prime}$, and a there is a homeomorphism of $Q$ that maps $-^\hat{h}$ to the concatenation of $h$ and $C^{\prime}_{c_{1},c_{2}}$.
@@
Define $U^{\prime}=L^{\prime}\cup C^{\prime}\cup R^{\prime}$. Let $f:[0,1]\rightarrow Q$ be a curve with $f(x)\notin C^{\prime}$ for any $x\in[0,1)$. By continuity of $f$ there is some $\beta\in(0,1)$ for which $f((\beta,1])\subset U^{\prime}$. We claim $f((\beta,1))\subset L^{\prime}$ or $f((\beta,1])\subset R^{\prime}$. If $\phi(f((\beta,1)))$ contains a point in $C^{\prime}\times(-\infty,0)$ and a point in $\phi^{-1}(C^{\prime}\times(0,\infty)$ then by continuity $\phi(f((\beta,1)))$ contains a point in $C^{\prime}\times\{0\}$ and hence $f((\beta,1))$ contains a point in $C^{\prime}$ which is a contradiction. Thus either $\phi(f((\beta,1)))\subset C^{\prime}\times(-\infty,0)$ or $\phi(f((\beta,1)))\subset C^{\prime}\times(0,\infty)$. If $\phi(f((\beta,1)))\subset C^{\prime}\times(-\infty,0)$, then $f((\beta,1))\subset L^{\prime}$. If $\phi(f((\beta,1)))\subset C^{\prime}\times(0,\infty)$, then $f((\beta,1))\subset R^{\prime}$.
@@
For $t$ close enough to $t_{0}$ the last line is positive. This implies that for some $t_{2}>t_{0}$, $\|f(t)-f(t_{0})\|^{2}$ is increasing on $[t_{0},t_{2}]$. Likewise for some $t_{3}<t_{0}$, $\|f(t)-f(t_{0})\|^{2}$ is decreasing on $[t_{3},t_{0}]$.
Let $r>0$ be such that $\|f(t_{0})-f(t)\|\geq 2r$ for all $t\in[0,1]\backslash\operatorname{frac}([t_{3},t_{2}])$, where $\operatorname{frac}(x)$ is the fractional part of $x$. Then $\|f(t_{0})-f(t_{2})\|,\|f(t_{0})-f(t_{3})\|\geq 2r$. Then since $\|f(t_{0})-f(t)\|$ is increasing on $[t_{0},t_{2}]$ there is exactly one $t_{4}\in[t_{0},t_{2}]$ with $\|f(t_{0})-f(t_{4})\|=r$. Likewise there is exactly one $t_{5}\in[t_{3},t_{0}]$ with $\|f(t_{0})-f(t_{5})\|=r$. So $f([t_{5},t_{4}])$ forms a simple curve in the closed ball $\bar{B}(f(t_{0}),r)$ with endpoints on the boundary and $f((t_{5},t_{4}))$ lying in the interior. It follows from the Jordan curve theorem that $B(f(t_{0}),r)\backslash f([0,1])=B(f(t_{0}),r)\backslash f((t_{5},t_{4}))$ contains exactly two connected components. ∎
@@
By 3 $\phi_{2}$ crosses $\phi_{3}$. Let $L_{\phi_{3}},R_{\phi_{3}}$ be the left and right of $\phi_{3}$ as in 12. Since $L_{\phi_{3}},R_{\phi_{3}}$ are connected, $L_{\phi_{3}},R_{\phi_{3}}$ belong to different regions of $Q\backslash\phi_{3}$. Since $\phi_{2}$ crosses $\phi_{3}$, there exists $t_{0}$ for which $\phi(t_{0})\in Q_{1}$ and $t_{1}$ for which $\phi(t_{1})\in Q_{2}$. Let $t_{2}=\operatorname{inf}\{a\in[0,1]:\ \exists b\in(a,1]\ \text{s.t.}\ \phi((a,b))\subset Q_{1}\}$, $t_{3}=\operatorname{sup}\{b\in[t_{2},1]:\ \text{s.t.}\phi((t_{2},b))\subset Q_{1}\}$, that is $[t_{2},t_{3}]$ is a maximal interval for which $\phi([t_{2},t_{3}])\subset Q_{1}$. It follows $\phi(t_{2}),\phi(t_{3})$ lie on the boundary of $Q_{1}$, that is on $\phi_{1}$. This implies $t_{2},t_{3}\in\{0,1\}$. Since $t_{2}<t_{3}$ $t_{2}=0$ and $t_{3}=1$. This implies that $\phi([0,1]))$ lies in $Q_{1}\cup\{\phi(0)\}$ contradicting that there exists $t_{1}$ for which $\phi(t_{1})\in Q_{2}$.
@@
(of 15) Let $f:[0,1]\rightarrow C^{\prime}$ be a paramaterization of $C^{\prime}$ and let $c=f^{-1}(c_{1})$. Note that by 14 the curve $\phi_{2}$ is homeomorphic to a curve $\phi_{3}$ that enters $\phi$ from the right and leaves $\phi$ from the left at the same point $f(c)$. By 19, $\phi_{3}$ does not divide $Q$ into 2 regions. Thus, $\phi_{2}$ does not divide $Q$ into 2 regions.

### v4

@@ 4 Solving the case of no facial cycles
([1] p.162) For a surface without boundary $Q$ of genus $g^{\prime}$, $Q^{\prime}$ obtained by 3 is a surface without boundary of genus at most $g^{\prime}-1$.
We apply the surgery of 3 to $g(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ to obtain a surface $Q^{\prime}$ of genus one less than $Q$. Let $N^{\prime}_{1},N^{\prime}_{2}$ denote the two cones glued to $Q^{\prime}$. We also apply the surgery of 3 to $g(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ to obtain a surface $\hat{Q}$ of genus one less than $Q$. Let $\hat{N}_{1},\hat{N}_{2}$ denote the two cones glued to $\hat{Q}$.
Lemma 4.
Let $G$ be a graph embedded on a surface $Q$ with no dicycles. Let $h$ be a non-facial curve of $Q\backslash G$. Let $Q^{\prime}$ be the surface obtained by applying the surgery of 3 to with respect to the curve $h$ and surface $Q$. There is a natural embedding of $G$ on $Q^{\prime}$ (by leaving each node of $G$ where it was in $Q$). Let $N_{1},N_{2}$ denote the two cones glued to $Q^{\prime}$ during the surgery process. Then each facial cycle of $G$ with respect to its embedding in $Q^{\prime}$ contains either $N_{1}$ or $N_{2}$ in its inside region.
Proof.
Let $C$ be a facial cycle of $G^{F}\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ with respect to its embedding in $Q^{\prime}$. If neither of the cones $N_{1},N_{2}$ are contained in the inside region of $C$, then $C$ is a facial cycle of $G$ with respect to its embedding in $Q$, which is a contradiction. ∎
Thus, any facial cycle of $G^{F}\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ contains either $N^{\prime}_{1}$ or $N^{\prime}_{2}$ in its inside region.
Now let $G_{1},G_{2},..,G_{l}$ be the strongly connected components of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$. Since any closed walk of of $G^{F}\backslash(R^{\prime}\cup K^{\prime+})$ is either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ each strongly connected component is either contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ or in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$. If $G_{i}$ is contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$, then there is a natural embedding of $G_{i}$ in $Q^{\prime}$ (obtained by leaving all nodes and edges where they are in the surgery for 3). Likewise, if $G_{i}$ is contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$, then there is a natural embedding of $G_{i}$ in $\hat{Q}$. Thus, for any $G_{i}$ contained in $G\backslash(P_{1}\cup C^{1}_{(v_{a},v_{i})})$ by Theorem 2, there is an 8-approximation for the problem of hitting the facial cycles of $G_{i}$ (with respect to the natural embedding in $Q^{\prime}$). Likewise, for any $G_{i}$ contained in $G\backslash(P_{2}\cup C^{1}_{(v_{j},v_{b})})$ there is an 8-approximation for the problem of hitting the facial cycles of $G_{i}$. Let $Z_{i}$ be a solution for the problem of hitting facial cycles of $G_{i}$ of cost at most $8OPT_{LP(G_{i})}$ as guaranteed by Theorem 2.
Then each $G_{i}\backslash Z_{i}$ is embedded in a surface of smaller genus with no facial cycles.
By induction, there are solutions $A_{i}$ to $G_{i}\backslash Z_{i}$ of cost $c_{g-1}OPT_{LP(G_{i}\backslash Z_{i})}$, where $c_{g}$ is the integrality gap of the DFVS LP for graphs of genus $g$.
Define $\hat{x}^{G_{i}\backslash Z_{i}}\in\mathbb{R}^{V(G_{i}\backslash Z_{i})}$ as $\hat{x}^{G_{i}}(v)=\hat{x}_{v}$, where $\hat{x}$ is as in the proof of 3. Since graphs $G_{i}$ are vertex disjoint, $G_{i}\backslash Z_{i}$ are vertex disjoint, so $\sum_{i=1}^{l}OPT_{LP(G_{i})}\leq\sum_{i=1}^{l}\sum_{v\in V(G_{i})}\hat{x}^{G}_{i}(v)\leq\sum_{v\in V(G)}\hat{x}=OPT_{LP(G)}$. Now $F\cup R^{\prime}\cup K^{\prime+}\cup(\cup_{i=1}^{l}A_{i})\cup(\cup_{i=1}^{l}Z_{i})$ is a DFVS of cost $(O(1)+c_{g-1})OPT_{LP(G)}=(O(1)+O(g-1))OPT_{LP(G)}=(O(g))OPT_{LP(G)}$.
Note that the argument in 1 is symmetric with respect to left and right and we may swap right and left to get the following result. Let $b^{\prime}_{i,1},b^{\prime}_{i,2},..,b^{\prime}_{i,l^{\prime}_{i}}$ be the in neighbours of $v_{i}$ such that each edge $b^{\prime}_{i,t}v_{i}$ reaches $v_{i}$ from the left and $d^{\prime}_{i,1},d^{\prime}_{i,2},..,d^{\prime}_{i,t^{\prime}_{i}}$ be the out neighbours of $v_{i}$ such that the edge $d^{\prime}_{i,t^{\prime}}v_{i}$ reaches $v_{i}$ from the right. Subdivide each edge $b^{\prime}_{j,t^{\prime}}v_{j}$ into a path $b^{\prime}_{j,t^{\prime}}b_{j,t^{\prime}}v_{j}$ and each edge $v_{i}d^{\prime}_{i,t}$ into a path $v_{i}d_{i,t}d^{\prime}_{i,t}$ and give the new vertices $d_{j,t^{\prime}},b_{i,t}$ infinite cost. There is a natural embedding of our new graph on our surface by placing each $b_{i,t}$ where the midpoint of the curve $g(v_{i}b^{\prime}_{j,t^{\prime}})$ was embedded and likewise for $d_{j,t^{\prime}}$. By abuse of notation, we continue to call our graph $G$ and define $\hat{x}_{b_{i,t}}=\hat{x}_{d_{j,t^{\prime}}}=0$ for all $b_{i,t}$, $d_{j,t^{\prime}}$.
Claim 4.
If $d_{\omega}(V_{\kappa_{-}},V_{\kappa_{+}})\leq\frac{1}{12}N$, then the LP gap of the natural LP ($\mbox{P}_{\text{DFVS}}$) for $G$ has integrality gap $O(1)$.
Proposition 9.
If there is a (undirected) path $P=q_{1},q_{2},..,q_{t}w_{i,t}$ from some $w_{i,t}\in W$ (resp $b_{i,t}\in B$) in $\operatorname{un}(K)\backslash C^{1}$, then there is a directed path from $q_{j}$ to $W$ (resp $B$) in $G^{F}\backslash(S\cup T\cup C^{1})$ for any $j=1,2,..,t$.
Proof.
Let $P=u_{i,r}q_{1},q_{2},..,q_{t}$ be a path in $\operatorname{un}(K)\backslash C^{1}$ from some $u_{i,r}\in U$ (resp $u_{i,r}\in D$). We prove by induction $t^{\prime}$ on that there is a directed path from $U$ to $q_{j}$ in $G^{F}\backslash(S\cup T\cup C^{1})$ for any $j=1,2,..,t^{\prime}$. The case $t^{\prime}=1$ is clear as each $u_{i,r}\in U$ (resp $u_{i,r}\in D$) only has out neighbours so the undirected edge $\{u_{i,r},q_{1}\}$ in $\operatorname{un}(K)$ is directed from $u_{i,r}$ to $q_{1}$.
Now assume the statement true for $t^{\prime}=t^{\prime\prime}$. For $t^{\prime}=t^{\prime\prime}+1$, if the undirected edge $\{q_{t^{\prime}},q_{t^{\prime}+1}\}$ is directed from $q_{t^{\prime}}$ to $q_{t^{\prime}+1}$, then there is a directed path from $u_{i,r}$ to $q_{t^{\prime}+1}$ in $G^{F}\backslash(S\cup T\cup C^{1})$.
Otherwise $\{q_{t^{\prime}},q_{t^{\prime}+1}\}$ is directed from $q_{t^{\prime}+1}$ to $q_{t^{\prime}}$. By strong connectedness of $K$, there is a directed path $P^{\prime}$ from $q_{t^{\prime}}$ to $q_{t^{\prime}+1}$ in $K\backslash(S\cup T)$. If $P^{\prime}$ does not intersect $C$ then there is a directed path from $u_{i,r}$ to $q_{t^{\prime}+1}$ in $G^{F}\backslash(S\cup T\cup C^{1})$. So, assume $P^{\prime}$ intersects $W$ or $B$. Let $P^{\prime\prime}$ denote the subpath of $P^{\prime}$ from $q_{t^{\prime}}$ to when $P^{\prime}$ first intersects $U$ or $B$. By construction $P^{\prime\prime}$ lies in $G^{F}\backslash(S\cup T\cup C^{1})$. As $u_{i,r}$ lies in $U$ (resp $D$) $P^{\prime\prime}$ does not intersect $W$ (resp. $B$), as then we would have a $U$-$W$ (resp. $D$-$B$) path in $G^{F}\backslash(S\cup T\cup C^{1})$. Thus, $P^{\prime\prime}$ is a $q_{t^{\prime}}$-$B$ (resp. $q_{t^{\prime}}$-$W$) path. Consider the subpath $Q$ of the reversal of $P^{\prime}$ starting from $q_{t^{\prime}+1}$ to when the reversal of $P^{\prime}$ first intersects $D$ or $U$. Let $\operatorname{rev}(Q)$ denote the reversal of $Q$. Note $\operatorname{rev}(Q)$ lies in $G^{F}\backslash(S\cup T\cup C^{1})$. If the starting vertex of $\operatorname{rev}(Q)$ is in $D$ (resp. $U$), then $\operatorname{rev}(Q)\cup\{q_{t^{\prime}+1}q_{t^{\prime}}\}\cup P^{\prime\prime}$ is a $D$-$B$ (resp. $U$-$W$) path in $G^{F}\backslash(S\cup T\cup C^{1})$. This contradicts 4. Thus, the starting vertex of of $\operatorname{rev}(Q)$ is in $U$ (resp. $D$). This implies there is a path from $U$ (resp. $D$) to $q_{t^{\prime}+1}$ completing the induction. The proof of the second part is similar. ∎
Proposition 10.
Proof.
If we have a $U$-$W$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $\operatorname{un}(K)\backslash C^{1}$, then by 9, there are directed $U$-$q_{1}$ and $q_{1}$-$W$ paths $P_{1}$ and $P_{2}$ in $\operatorname{un}(K)\backslash C^{1}$. Then $P_{1}\cup P_{2}$ is a directed $U$-$W$ path in $K\backslash C^{1}$ which contradicts 4. Thus, we do not have a $U$-$W$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $K\backslash C^{1}$. Likewise, we do not have a $D$-$B$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}w_{j,t^{\prime}}$ in $K\backslash C^{1}$.
Suppose we have a $U$-$D$ path $P=u_{i,t}q_{1},q_{2},..,q_{t}d_{j,t^{\prime}}$ in $\operatorname{un}(K)\backslash C^{1}$. By 9, there are directed $U$-$q_{1}$ and $D$-$q_{1}$ paths $P_{1}$ and $P_{2}$ in $K\backslash C^{1}$. Recall $U$ has no in-neighbours of in $G\backslash C^{1}$, so the edge $\{u_{i,t},q_{1}\}$ in $K$ is directed from $u_{i,t}$ to $q_{1}$. By 2 connectedness of $K$, there is a path $P_{3}$ from $q_{1}$ to $u_{i,t}$. The only in-neighbours of $u_{i,t}$ are in $C^{1}$, thus $P_{3}$ intersects $W\cup B$. Let $P^{\prime}_{3}$ be the subpath of $P_{3}$ from $q_{1}$ to when it the path first intersects $W\cup B$. If the endpoint of $P^{\prime}_{3}$ is in $W$, then $P_{1}\cup P^{\prime}_{3}$ is a $U$-$W$ path in $K\backslash C^{1}$. Otherwise, if the endpoint of $P^{\prime}_{3}$ is in $B$, then $P_{2}\cup P^{\prime}_{3}$ is a $D$-$B$ path in $K\backslash C^{1}$. Either way this contradicts 4.
Proposition 11.
[15][25]Suppose $G$ is a graph embedded on a surface $Q$. Let $C$ be a cycle of $G$ that does not divide $Q$ into two separate regions such that there is no edge between vertices of $C$ that is not part of $C$. Define a “left” and “right” as in 6. Let $\hat{L}$, $\hat{R}$ denote the neighbours of $C$ that are “left” or “right” of $C$. Suppose each connected component of $G\backslash C$ only contains nodes of $\hat{L}$ or $\hat{R}$ but not both. There is a non-facial closed curve $h$ in $Q\backslash G$.
Applying 10, we get that $G^{F}$ satisfies 11 with respect to $C^{1}$. Thus, there is a non-facial closed curve $h$ in $Q\backslash G^{F}$. We apply the surgery of 3 with respect to the closed curve $h$ and surface $Q$ to obtain a surface $Q^{\prime}$ of lower genus. Let $G_{1},G_{2},.,G_{l}$ be the strongly connected components of $G^{F}\backslash(S\cup T)$, so each $G_{i}$ is embeddable on $Q^{\prime}$. By 4 each facial dicycle of $G_{i}$ contains one of the cones of $Q^{\prime}$. Hence there is an algorithm that returns a hitting set $Z_{i}$ to the set of facial cycles of $G_{i}$ of cost at most $8OPT_{LP(G_{i})}$. By induction, there are solutions $A_{i}$ to $G_{i}\backslash Z_{i}$ of cost $c_{g-1}OPT_{LP(G_{i})}$, where $c_{g}$ is the integrality gap of the DFVS LP for graphs of genus $g$. Define $\hat{x}^{G}_{i}\in\mathbb{R}^{V(G_{i})}$ as $\hat{x}^{G}_{i}(v)=\hat{x}_{v}$, where $\hat{x}$ is as in the proof of 3. Since graphs $G_{i}$ are vertex disjoint, $\sum_{i=1}^{l}OPT_{LP(G_{i})}\leq\sum_{i=1}^{l}\sum_{v\in V(G_{i})}\hat{x}^{G}_{i}(v)\leq\sum_{v\in V(G)}\hat{x}=OPT_{LP(G)}$. Then $S\cup T\cup F\cup(\cup_{i=1}^{l}A_{i})\cup(\cup_{i=1}^{l}Z_{i})$ is a DFVS of cost $(O(1)+c_{g-1})OPT_{LP(G)}=(O(1)+O(g-1))OPT_{LP(G)}=(O(g))OPT_{LP(G)}$.
As observed in [7], ($\mbox{P}_{\text{DFVS}}$) can be solved in polynomial-time via the ellipsoid method. Hence 3 yields a polynomial time $O(g)$-approximation algorithm for DFVS in graphs of genus $g$ with no facial cycle.
5 Statement and proofs of topological results we use
First let us prove 11.
Proof.
@@ 5 Statement and proofs of topological results we use
Suppose, without loss of generality, that $G_{L}$ is empty. Consider the face of $G$ that contains $C$ and intersects the left of $C$. But, $C$ is not contractible (else it would separate the surface $Q$ into two components). Hence, a small leftward shift of $C$ which will lie in the face $f$ will produce a non-facial curve $h$.
@@
We claim that if a face contains vertices of $G_{L},G_{R}$ and of $C$ then there is a non-facial curve in $Q\backslash G$. Let $f$ be such a face of degree $d$. Let $\partial f=v_{0}\dotsm v_{d-1}$ be the boundary cycle of $f$, where $i\in\mathbb{Z}/d\mathbb{Z}$. Without loss of generality, assume that $v_{0}\in\hat{L}$ and for some $q$ $v_{1},v_{2},..,v_{q}\in C$, and $v_{q+1}\in\hat{R}$. There is are points $p_{l}$ on the edge $v_{0}v_{1}$ in the interior of $L$ and $p_{R}$ on the edge $v_{q}v_{q+r}$ in the interior of $R$. Let $h:[0,1]\rightarrow Q$ be a non-self-intersecting curve in $f$ from $p_{R}$ to $p_{L}$. Let $r_{L},r_{R}>0$ be such that $B_{Q}(p_{l},r_{L})\subset L,\ \ B_{Q}(p_{R},r_{R})\subset R$. Let $h^{L},h^{R}$ be non-self-intersecting curves in $B_{Q}(p_{l},r_{L})$ and $B_{Q}(p_{R},r_{R})$ from $p_{l}$ to $v_{1}$ and $p_{R}$ to $v_{q}$ respectively not intersecting $h$. Then $h\cup h^{L}\cup h^{R}$ satisfies the conditions of 7. Thus, $h\cup h^{L}\cup h^{R}\cup g(p_{l}v_{0},v_{1},...,v_{q+1}p_{R})$ does not bound a region of the closure of $f$. As this curve lies in the closure of $f$, this implies that $f$ is not homeomorphic to an open disk. By the classification theorem for orientable surfaces (see for instance page 87 of [17] ), $\operatorname{cl}(f)$ is homeomorphic to a $m$-torus $T_{m}$ with a finite number of open disks removed. Since $f$ is not homeomorphic to an open disk, $f$ contains a non-facial closed curve $h$ in its interior.
If there is a face $f$ of $G$ whose boundary contains vertices of $G_{L}$ and of $G_{R}$ (but not of $C$), then as there is no edge between $G_{L}$ and $G_{R}$, the boundary of $f$ is not connected and so $f$ is not homeomorphic to an open disk. Just as before this implies $f$ contains a non-facial closed curve $h$.
@@
It is well known (see for instance [1] page 15) that smooth surfaces $Q$ have the property that for each $v\in Q$ there is an open ball $B_{Q}(v,r_{0})$ of some small radius $r_{0}>0$ in $Q$ and a diffeomorphism $\psi$ from $B_{Q}(v,r_{0})$ to the open disk $B_{\mathbb{R}^{2}}(0,r_{0})$ of radius $r_{0}$ about the origin in the two-dimensional plane such that $\psi(v)=(0,0)$ and $\psi$ preserves distances from $v$, that is $\operatorname{dist}_{Q}(v,x)=\|\psi(v)-\psi(x)\|$, where $\operatorname{dist}_{Q}(v,x)$ is the geodesic distance from $v$ to $x$ in $Q$. For $p\in Q$, $r>0$, denote by $B(p,r)$ the open ball of radius $r$ about $p$. We now formally state and prove what 6 and 7 informally say.
@@
Lastly let $h:[0,1]\rightarrow Q$ be any curve that reaches $C^{\prime}$ from the right at a point $c_{2}=h(1)$ on $C^{\prime}$, leaves $C^{\prime}$ from the left at $c_{1}=h(0)$, that is the curve $\bar{\psi}(t)=h(1-t)$ reaches $C^{\prime}$ at $c_{1}$ from the left and $h$ is otherwise disjoint from $C^{\prime}$. Assume $c_{1}\neq c_{2}$ and let $C^{\prime}_{c_{1},c_{2}}$ be a subcurve of $C^{\prime}$ with endpoints $c_{1}$ and $c_{2}$. Then there is a curve $\hat{h}:[0,2]\rightarrow Q$ that reaches $C^{\prime}$ from the right at $c_{1}=\hat{h}(1)$ and leaves $C^{\prime}$ at a point $c_{1}=\hat{h}(0)$ $\hat{h}$ is otherwise disjoint from $C^{\prime}$, and a there is a homeomorphism of $Q$ that maps $\hat{h}$ to the concatenation of $h$ and $C^{\prime}_{c_{1},c_{2}}$.
@@
Define $U^{\prime}=L^{\prime}\cup C^{\prime}\cup R^{\prime}$. Let $f:[0,1]\rightarrow Q$ be a curve with $f(x)\notin C^{\prime}$ for any $x\in[0,1)$. By continuity of $f$ there is some $\beta\in(0,1)$ for which $f((\beta,1])\subset U^{\prime}$. We claim $f((\beta,1))\subset L^{\prime}$ or $f((\beta,1])\subset R^{\prime}$. If $\phi(f((\beta,1)))$ contains a point in $C^{\prime}\times(-\infty,0)$ and a point in $\phi^{-1}(C^{\prime}\times(0,\infty)$ then by continuity $\phi(f((\beta,1)))$ contains a point in $C^{\prime}\times\{0\}$ and hence $f((\beta,1))$ contains a point in $C^{\prime}$ which is a contradiction. Thus, either $\phi(f((\beta,1)))\subset C^{\prime}\times(-\infty,0)$ or $\phi(f((\beta,1)))\subset C^{\prime}\times(0,\infty)$. If $\phi(f((\beta,1)))\subset C^{\prime}\times(-\infty,0)$, then $f((\beta,1))\subset L^{\prime}$. If $\phi(f((\beta,1)))\subset C^{\prime}\times(0,\infty)$, then $f((\beta,1))\subset R^{\prime}$.
@@
For $t$ close enough to $t_{0}$ the last line is positive. This implies that for some $t_{2}>t_{0}$, $\|f(t)-f(t_{0})\|^{2}$ is increasing on $[t_{0},t_{2}]$. Likewise, for some $t_{3}<t_{0}$, $\|f(t)-f(t_{0})\|^{2}$ is decreasing on $[t_{3},t_{0}]$.
Let $r>0$ be such that $\|f(t_{0})-f(t)\|\geq 2r$ for all $t\in[0,1]\backslash\operatorname{frac}([t_{3},t_{2}])$, where $\operatorname{frac}(x)$ is the fractional part of $x$. Then $\|f(t_{0})-f(t_{2})\|,\|f(t_{0})-f(t_{3})\|\geq 2r$. Then since $\|f(t_{0})-f(t)\|$ is increasing on $[t_{0},t_{2}]$ there is exactly one $t_{4}\in[t_{0},t_{2}]$ with $\|f(t_{0})-f(t_{4})\|=r$. Likewise, there is exactly one $t_{5}\in[t_{3},t_{0}]$ with $\|f(t_{0})-f(t_{5})\|=r$. So $f([t_{5},t_{4}])$ forms a simple curve in the closed ball $\bar{B}(f(t_{0}),r)$ with endpoints on the boundary and $f((t_{5},t_{4}))$ lying in the interior. It follows from the Jordan curve theorem that $B(f(t_{0}),r)\backslash f([0,1])=B(f(t_{0}),r)\backslash f((t_{5},t_{4}))$ contains exactly two connected components. ∎
@@
By 5 $\phi_{2}$ crosses $\phi_{3}$. Let $L_{\phi_{3}},R_{\phi_{3}}$ be the left and right of $\phi_{3}$ as in 12. Since $L_{\phi_{3}},R_{\phi_{3}}$ are connected, $L_{\phi_{3}},R_{\phi_{3}}$ belong to different regions of $Q\backslash\phi_{3}$. Since $\phi_{2}$ crosses $\phi_{3}$, there exists $t_{0}$ for which $\phi(t_{0})\in Q_{1}$ and $t_{1}$ for which $\phi(t_{1})\in Q_{2}$. Let $t_{2}=\operatorname{inf}\{a\in[0,1]:\ \exists b\in(a,1]\ \text{s.t.}\ \phi((a,b))\subset Q_{1}\}$, $t_{3}=\operatorname{sup}\{b\in[t_{2},1]:\ \text{s.t.}\phi((t_{2},b))\subset Q_{1}\}$, that is $[t_{2},t_{3}]$ is a maximal interval for which $\phi([t_{2},t_{3}])\subset Q_{1}$. It follows $\phi(t_{2}),\phi(t_{3})$ lie on the boundary of $Q_{1}$, that is on $\phi_{1}$. This implies $t_{2},t_{3}\in\{0,1\}$. Since $t_{2}<t_{3}$ $t_{2}=0$ and $t_{3}=1$. This implies that $\phi([0,1]))$ lies in $Q_{1}\cup\{\phi(0)\}$ contradicting that there exists $t_{1}$ for which $\phi(t_{1})\in Q_{2}$.
@@
(of 15) Let $f:[0,1]\rightarrow C^{\prime}$ be a parameterization of $C^{\prime}$ and let $c=f^{-1}(c_{1})$. Note that by 14 the curve $\phi_{2}$ is homeomorphic to a curve $\phi_{3}$ that enters $\phi$ from the right and leaves $\phi$ from the left at the same point $f(c)$. By 19, $\phi_{3}$ does not divide $Q$ into 2 regions. Thus, $\phi_{2}$ does not divide $Q$ into 2 regions.